4. (15 points) A solution that is 0.500 M HF and 0.240 M NaOH reacts to produce a mixture with what pH? The pKa of HF is 3.20(15 points) A substance MX has a Ksp of 1.27 x 10-36. a. What is its molar solubility in pure water? (hint: ICE Table-remind- What is its molar solubility in 0.25 MBr? (hint: ICE Table-but what is M2 initial?) Please explain, thank you.
Answer #1

0.24 co -H F + Naon Na F +4₂0 6.50 0.24 This is a acidic Orso -0.24 o पिfer Sector the strong Base is not present in the soilqh = pk + dag skall} pt 3.20 + logo2.14 0.24 0.5-0.14 = 3,20 + Tog 0:24 0.26 PM = 3.20 + lsyo.92 P = 3.200 0.03 [ph = 3.17)To 27 XI-36, because neglect the - без релл •gs] 1.27 x 1036 - sxo.25- SESE 127 X1036 0. 25 s e 5.08x10-36 M